OpenStudy (anonymous):

How much water must be added to 12 L of a 40% solution of alcohol to obtain a 30% solution? A. .5 L B. 5 L C. .4 L D. 4 L

OpenStudy (anonymous):

almost identical to the last one

OpenStudy (mathmate):

The total volume should be 12l*(40%/30%)=16l So how much water must be added?

Parth (parthkohli):

@emmat9795 Please close a problem only when you finish with it. Thanks :)

OpenStudy (anonymous):

you have \(.4\times 12=4.8\) L of alcohol to begin with if you add say \(x\) liters of water you will still have 4.8 L of alcohol but now you will have \(12+x\) liters of liquid you want this to be 30% alcohol, so set \(4.8=0.3(12+x)\) and solve for \(x\)

OpenStudy (anonymous):

whatever @ParthKohli